# Universal Modelling of Emergent Oscillations in Fractional Quantum Hall Fluids

### v1

@@ Abstract
Density oscillations in quantum fluids can reveal their fundamental characteristic features. In this work, we study the density oscillation of incompressible fractional quantum Hall (FQH) fluids created by flux insertion. For the model Laughlin state, we find that the complex oscillations seen in various density profiles in real space can be universally captured by a simple damped oscillator model in the occupation-number space. It requires only two independent fitting parameters or characteristic length scales: the decay length and the oscillation wave number. Realistic Coulomb quasiholes can be viewed as Laughlin quasiholes dressed by magnetorotons which can be modeled by a generalized damped oscillator model. Our work reveals the fundamental connections between the oscillations seen in various aspects of FQH fluids such as in the density of quasiholes, edge, and the pair correlation function. The presented model is useful for the study of quasihole sizes for their control and braiding in experiments and large-scale numerical computation of variational energies.
@@
Introduction.— Characteristic density oscillations in the presence of perturbations are a fundamental aspect of quantum fluids. For the Fermi liquid, the long-range Friedel oscillation in the presence of impurity is a direct consequence of the existence of Fermi surface and quasiparticle excitations [1]. Similarly, for the non-Fermi liquids which go beyond the Fermi liquid paradigm (e.g. the Luttinger liquid [2], composite fermion liquid [3], quantum Hall liquid [4, 5] and strange metal [6]), characteristic density oscillations manifest as spin-charge separation [7], charge-vortex duality [8], anomalous decay laws and exponents [9, 10, 11, 12, 13], and so on. Understanding and accurately modeling such oscillations, which are of both theoretical and experimental significance, remains an outstanding open problem.
@@
An important class of non-Fermi liquids is the incompressible fractional quantum Hall (FQH) fluids which display characteristic oscillatory features that encode both geometric and universal topological information. When a quantum of flux is inserted into the uniform FQH ground state, quasiholes carrying a fractional charge [14] and obeying fractional statistics [15] are created. They also carry a dipole moment to balance the Hall viscosity in the presence of the electric field gradient, which is proportional to the FQH topological shift [16]. The dipole moment, a characteristic feature of incompressibility, is established from the density oscillation at the edge [17]. Moreover, when an appropriate number of fluxes equivalent to removing an electron is inserted at the same position, the density of the bound state of a few stacked quasiholes is proportional to the pair-correlation function of the ground state [18]. In the limit of an infinite number of fluxes inserted, a macroscopic FQH edge is created, near which the density oscillation has been intensively studied [19, 18, 20, 21, 17, 22, 23, 24, 25, 26]. The FQH quasiholes, ground state pair-correlation function, and the QH edge can thus be understood as special cases of flux insertion (see Fig. 1) and are thus closely related. The precise underlying connections and both the qualitative and quantitative aspects of such oscillations, however, are not well understood.
@@
In this Letter, we show that the real space density oscillation from the flux insertion in the model Laughlin state can be accurately modeled by a simple damped oscillation with degrees of freedom within a single Landau level (LL). The oscillation is determined by two characteristic length scales: the decay length and the oscillation wave number, and these serve as the only fitting parameters of the model. In contrast to previous works that directly model the real space density [27, 20, 21, 28, 22, 11, 12, 25, 29], we emphasize that the model should only focus on the guiding center degrees of freedom (within a single LL), as those are the relevant coordinates for any FQH phase. For the model Laughlin state, we study its quasihole, edge, and pair-correlation function using the aforementioned model-based approach treating them all on an equal footing. A phenomenological model for the damped oscillation of Laughlin quasiholes is proposed at general fillings. For the more realistic Coulomb interaction, where quasiholes are dressed by neutral excitations, a generalized damping model with four characteristic lengths is shown to work very well, which is useful for both numerical computations and experimental manipulation of quasiholes. Moreover, we find this generalized model also gives an accurate description of quasiholes in some of the composite fermion and non-Abelian topological phases.
@@
Universal damped oscillation.— The model wave function of the Laughlin quasihole located at the origin of the disk geometry for $N_{e}$ electrons is $\Psi_{L}^{1{-}\text{qh}}{=}\left(\prod_{i{=}1}^{N_{e}}z_{i}\right)\Psi_{L},$ where $\Psi_{L}{=}\prod_{1{\leq}i{<}j{\leq}N_{e}}(z_{i}{-}z_{j})^{m}e^{-\frac{1}{4}\sum_{i{=}1}^{N_{e}}|z_{i}|^{2}}$ is the Laughlin wave function at $\nu{=}1/m$[14], $z_{i}{=}x_{i}{-}iy_{i}$ is the position of the $i^{\text{th}}$ electron, and the magnetic length at magnetic field $B$ is taken as the unit of length i.e., we set $l_{B}{=}\sqrt{\hbar c/(eB)}{=}1$. We can also stereographically map this state to the spherical geometry [30]. The density distribution of a single Laughlin quasihole at $\nu{=}1/3$ in the spherical geometry is shown in Fig. 2(a). Directly finding a simple empirical model for the real-space density distribution of quasiholes is challenging as the real-space structure is a mixture of the trivial LL (cyclotron) and the non-trivial FQH (guiding-center) contributions. A real-space density distribution $\rho(r)$ can be decomposed as $\rho(r){=}\sum_{i{=}0}^{N_{\phi}}n_{i}\rho_{i}(r)$, where $n_{i}$ is the average occupation number of the $i^{\text{th}}$ orbital, $\rho_{i}$ is the density computed from single-particle wave functions, and $N_{\phi}$ is the number of fluxes threading the sample. Only $n_{i}$ is related to the correlated FQH physics.
@@
where $k_{1}$ and $\lambda_{1}$ are the oscillation wave number and decay length respectively, $A_{1}$ is the amplitude and $x_{1}$ is the zero point. The fitted result (red crosses) agrees almost perfectly [31] with the exact occupations (blue circles) as shown in Fig. 2(b). Correspondingly, the exact and fitted real-space density distributions are also nearly indistinguishable from each other as shown in Fig. 2(a).
Figure 2: (a) The density distribution of a Laughlin quasihole located at the north pole of a sphere as a function of the arc distance $r$. The blue line and crosses denote the exact distribution and the fitting with our damped model [Eq. (1)], respectively for $N_{e}{=}17$. The red line and crosses denote the thermodynamic limit obtained from the polynomial expansion method [29] and our fitting [31], respectively, and the data points are shifted down by $0.01$ for clarity. (b) The occupation-number density. Blue circles and red crosses denote the exact density and that obtained from fitting with our model, respectively. The fitted curve is $n(r){=}0.68\sin[1.43(r{-}2.31)]\exp({-}r/1.21){+}\bar{n}$ with $\bar{n}{=}(N_{e}{+}1/3)/N_{o}$ with $N_{o}{=}50$.
It is important to note that the accurate fitting is achieved with only two independent fitting parameters $k_{1}$ and $\lambda_{1}$. The other two parameters in Eq. (1), $A_{1}$ and $x_{1}$, are fixed by the total charge $\sum_{i{=}0}^{N_{\phi}}\nu_{i}{=}N_{e}$ and the total angular momentum $\sum_{i{=}0}^{N_{\phi}}\nu_{i}(i{-}N_{\phi}/2){=}N_{e}/2$. The two conditions are equivalent to the constraints that a single quasihole has a charge $e_{1{-}\text{qh}}{=}{-}1/m$ and a dipole moment $d_{1{-}\text{qh}}{=}(1{-}1/m)/2$ in the thermodynamic limit; with both being topological and thus robust against perturbations. From finite-size-scaling, we find the following values of the parameters in the thermodynamic limit [31]:
@@
Stacked Laughlin quasiholes.— The damped oscillation model can also be applied to $n$-stacked quasiholes created by inserting $n$ fluxes at the same location. The real-space density of such quasiholes gets increasingly complex with increasing $n$. Even for IQH fluid, the $n$-stacked holes have $\delta\rho(r){=}{-}1/(2\pi)\exp\left({-}r^{2}/2\right)(\sum_{i{=}0}^{n{-}1}r^{2i}/2^{i}i!)$, which is no longer a simple Gaussian distribution for $n{>}1$. Focusing on the FQH system with multiple quasiholes, we again look at its occupation-number density which reveals interesting physics and is no more complex even for $n{>}1$. The values of the characteristic lengths $k_{1}^{(n)}$ and $\lambda_{1}^{(n)}$ for different values of $n$ are shown in Figs. 3(a) and 3(b), where both $k_{1}^{(n)}$ and $\lambda_{1}^{(n)}$ increase as $n$ gets larger. The complicated real-space densities can be easily restored with the single-particle wave functions.
@@
The $m$-stacked quasihole at $\nu{=}1/m$ deserves special attention because its density distribution is proportional to the pair-correlation function of the Laughlin ground state. The density distribution of $m$-Laughlin quasiholes is $\rho(r){=}\mathcal{N}_{1}\int\prod_{i{=}2}^{N_{e}^{\prime}}d\bm{r}_{i}|\Psi_{L}^{m{-}\text{qh}}(\bm{r},\bm{r}_{2},{\ldots},\bm{r}_{N_{e}^{\prime}})|^{2},$ where $\mathcal{N}_{1}$ is a constant. On the other hand, the pair-correlation function $g(r){=}\mathcal{N}_{2}\int\prod_{i{=}3}^{N_{e}}d\bm{r}_{i}|\Psi_{L}(0,\bm{r},\bm{r}_{3},{\ldots},\bm{r}_{N_{e}})|^{2}.$ By substituting the explicit wave function $\Psi_{L}$ and $\Psi_{L}^{m{-}\text{qh}}$ into the expression of $g(r)$ and $\rho(r)$, respectively, one can show that $\rho(r)$ is proportional to $g(r)$ when $N_{e}{=}N_{e}^{\prime}{+}1$. Therefore, the decay length of the density of $m$-stacked Laughlin quasiholes is equal to the correlation length $\lambda_{\text{cor}}$ between electrons:
@@
Our result is thus useful for the large-scale numerical computation of ground-state variational energies. One can use our modeled $g(r)$ to calculate the per-particle variational energy of the Laughlin state in terms of any general interaction $V(r)$ through the formula $V{=}(\bar{\rho}/2)\int d^{2}\,\bm{r}V(r)[g(r){-}1]$[32], where $\bar{\rho}$ is the average density. For the Coulomb interaction $V(r){=}1/r$, the calculated energy $V$ with our modeled $g(r)$ is $-0.4096$ for $\nu{=}1/3$ and $-0.3278$ for $\nu{=}1/5$[31], which are very close to the conjectured thermodynamic values $-0.4098$[33, 34] and $-0.3275(1)$[35] obtained from the extrapolation of small-size exact diagonalization results and large-scale Monte Carlo calculations. Thus for other realistic interactions, the thermodynamic variational energies can now be computed very efficiently with our modeled $g(r)$[31]. Besides, our method does not suffer from the systemic error of the polynomial expansion method at small $r$[29].
In the limit of $n{\rightarrow}\infty$, the corresponding “Laughlin quasihole” also deserves special attention as it represents the edge of the QH fluid. Its density profile can also be well-fitted with a damped model in the occupation-number space with the characteristic lengths being $k_{1}^{\text{edge}}{=}1.49$ and $\lambda_{1}^{\text{edge}}{=}1.51$[31] which are just the limiting values of $n$-quasiholes as shown in Fig. 3.
@@
Given the near-perfect fitting of such a simple model with only two fitting parameters [31], we conjecture that the damped oscillation model captures the main features of the universal oscillations in the Laughlin FQH fluid, and in principle can be derived. While we are not able to accomplish that here, a phenomenological model for Laughlin quasiholes at general fillings can be proposed based on these results. We view the system as a damped oscillator with wave number $k^{\prime}_{1}{=}\sqrt{k_{1}^{2}{+}\lambda_{1}^{-2}}$ serving as the intrinsic “wave number” which is a little larger than the wave number of the Wigner crystal $k_{\text{Wigner}}$ for small $m$ [see Fig. 3(c)]. The damping results from a “frictional force” which is proportional to the density $\nu{=}1/m$. The differential equation that describes this system is
@@
where $\mathcal{O}(\delta n^{2})$ denote terms that can potentially arise from non-linear effects which we ignore here. The linear solution of Eq. (4) is just our damped model given in Eq. (1). Note that the model proposed in Eq. (4) is different from that proposed in Eq. (23) of Ref. [21], which does not consider the damped term and assumes a different intrinsic wave number at the level of linear response. Deriving this simple model from the microscopic wave function and capturing the non-linear effects remains an open problem [21, 22].
Quasiholes for realistic interactions.— Going beyond the model Laughlin quasihole wave functions, the density profile of quasiholes from realistic interactions becomes more complicated, because such quasiholes are dressed by neutral excitations [37, 38]. Here we study the quasihole in the presence of the following tunable interaction

### v2

# Universal Modelling of Oscillations in Fractional Quantum Hall Fluids

@@ Abstract
Density oscillations in quantum fluids can reveal their fundamental characteristic features. In this work, we study the density oscillation of incompressible fractional quantum Hall (FQH) fluids created by flux insertion. For the model Laughlin state, we find that the complex oscillations seen in various density profiles in real space can be universally captured by a simple damped oscillator model in the occupation-number space. It requires only two independent fitting parameters or characteristic length scales: the decay length and the oscillation wave number. Realistic Coulomb quasiholes can be viewed as Laughlin quasiholes dressed by magnetorotons, which can be modeled by a generalized damped oscillator model. Our work reveals the fundamental connections between the oscillations seen in various aspects of FQH fluids such as in the density of quasiholes, edge, and the pair correlation function. The presented model is useful for the study of quasihole sizes for their control and braiding in experiments and large-scale numerical computation of variational energies.
I Introduction
@@
Characteristic density oscillations in the presence of perturbations are a fundamental aspect of quantum fluids. For the Fermi liquid, the long-range Friedel oscillation in the presence of impurity is a direct consequence of the existence of the Fermi surface and quasiparticle excitations [1]. Similarly, for the non-Fermi liquids and strongly interacting topological fluids which go beyond the Fermi liquid paradigm (e.g. the Luttinger liquid [2], composite fermion liquid [3], quantum Hall liquid [4, 5] and strange metal [6]), characteristic density oscillations manifest as spin-charge separation [7], charge-vortex duality [8], anomalous decay laws and exponents [9, 10, 11, 12, 13], and so on. Understanding and accurately modeling such oscillations, which are of both theoretical and experimental significance, remains an outstanding open problem.
@@
An important class of strongly interacting topological fluids is the incompressible fractional quantum Hall (FQH) fluids which display characteristic oscillatory features that encode both geometric and universal topological information. When a quantum of flux is inserted into the uniform FQH ground state, quasiholes carrying a fractional charge [14] and obeying fractional statistics [15] are created. They also carry a dipole moment to balance the Hall viscosity in the presence of the electric field gradient, which is proportional to the FQH topological shift [16]. The dipole moment, a characteristic feature of incompressibility, is established from the density oscillation at the edge [17]. Moreover, when an appropriate number of fluxes equivalent to removing an electron is inserted at the same position, the density of the bound state of the stacked quasiholes is proportional to the pair-correlation function of the ground state [18]. In the limit of an infinite number of fluxes inserted, a macroscopic FQH edge is created, near which the density oscillation has been intensively studied [19, 18, 20, 21, 17, 22, 23, 24, 25, 26]. The FQH quasiholes, ground state pair-correlation function, and the edge can be understood as special cases of flux insertion (see Fig. 1) and are thus closely related. The precise underlying connections and both the qualitative and quantitative aspects of such oscillations, however, are not well understood.
@@
In this paper, we show that the real space density oscillation from the flux insertion in the model Laughlin state can be accurately modeled by a simple damped oscillation with degrees of freedom within a single Landau level (LL). The oscillation is determined by two characteristic length scales: the decay length and the oscillation wave number, and these serve as the only fitting parameters of the model. In contrast to previous works that directly model the real space density [27, 20, 21, 28, 22, 11, 12, 25, 29], we emphasize that the model should only focus on the guiding center degrees of freedom (within a single LL), as those are the relevant coordinates for any FQH phase. For the model Laughlin state, we study its quasihole, edge, and pair-correlation function using the aforementioned model-based approach treating them all on an equal footing. A phenomenological model for the damped oscillation of Laughlin quasiholes is proposed at general fillings. For the more realistic Coulomb interaction, where quasiholes are dressed by neutral excitations, a generalized damping model with four characteristic lengths is shown to work very well. This general model can be useful for both numerical computations and experimental manipulation of quasiholes. Moreover, we find this generalized model also gives an accurate description of quasiholes in some of the composite fermion and non-Abelian topological phases.
The remainder of the paper is organized as follows. In Sec. II, we study the density oscillation of a single Laughlin quasihole. In Sec. III, we generalize the study to various stacked Laughlin quasiholes. Moving forward, we examine the density oscillation of quasihole for realistic interactions in Sec. IV, and for Jain and non-Abelian phases in Sec. V. Thereafter, we also investigate the oscillatory behaviors for quasielectrons in Sec. VI. Finally, we summarize our results in Sec. VII.
@@
II Universal damped oscillation
The model wave function of the Laughlin quasihole located at the origin of the disk geometry for $N_{e}$ electrons is $\Psi_{\rm L}^{1{-}\text{qh}}{=}\left(\prod_{i{=}1}^{N_{e}}z_{i}\right)\Psi_{\rm L},$ where $\Psi_{\rm L}{=}\prod_{1{\leq}i{<}j{\leq}N_{e}}(z_{i}{-}z_{j})^{m}e^{-\frac{1}{4}\sum_{i{=}1}^{N_{e}}|z_{i}|^{2}}$ is the Laughlin wave function at $\nu{=}1/m$ [14], $z_{i}{=}x_{i}{-}iy_{i}$ is the position of the $i^{\text{th}}$ electron, and the magnetic length at magnetic field $B$ is taken as the unit of length i.e., we set $l_{B}{=}\sqrt{\hbar c/(eB)}{=}1$. We can also stereographically map this state to the spherical geometry [30]. The density distribution of a single Laughlin quasihole at $\nu{=}1/3$ in the spherical geometry is shown in Fig. 2(a). Directly finding a simple empirical model for the real-space density distribution of quasiholes is challenging as the real-space structure is a mixture of the trivial LL (cyclotron) and the non-trivial FQH (guiding-center) contributions. A real-space density distribution $\rho(r)$ can be decomposed as $\rho(r){=}\sum_{i{=}0}^{N_{\phi}}n_{i}\rho_{i}(r)$, where $n_{i}$ is the average occupation number of the $i^{\text{th}}$ orbital, $\rho_{i}$ is the density computed from single-particle wave functions, and $N_{\phi}$ is the number of fluxes threading the sample. Only $n_{i}$ is related to the correlated FQH physics.
@@ II Universal damped oscillation
where $k_{1}$ and $\lambda_{1}$ are the oscillation wave number and decay length respectively, $A_{1}$ is the amplitude and $x_{1}$ is the zero point. The fitted result (red crosses) agrees almost perfectly with the exact occupations (blue circles) as shown in Fig. 2(b). Correspondingly, the exact and fitted real-space density distributions are also nearly indistinguishable from each other as shown in Fig. 2(a).
Figure 2: (a) The density distribution of a Laughlin quasihole located at the north pole of a sphere as a function of the arc distance $r$. The blue line and crosses denote the exact distribution and the fitting with our damped model [Eq. (6)], respectively for $N_{e}{=}17$. The red line and crosses denote the thermodynamic limit obtained from the polynomial expansion method [29] and our fitting [see Appendix. B for details], respectively, and the data points are shifted down by $0.01$ for clarity. (b) The occupation-number density. Blue circles and red crosses denote the exact density and that obtained from fitting with our model, respectively. The fitted curve is $n(r){=}0.68\sin[1.43(r{-}2.31)]\exp({-}r/1.21){+}\bar{n}$ with $\bar{n}{=}(N_{e}{+}1/3)/N_{o}$ with $N_{o}{=}50$.
It is important to note that the accurate fitting is achieved with only two independent fitting parameters $k_{1}$ and $\lambda_{1}$. The other two parameters in Eq. (6), $A_{1}$ and $x_{1}$, are fixed by the total charge $\sum_{i{=}0}^{N_{\phi}}\nu_{i}{=}N_{e}$ and the total angular momentum $\sum_{i{=}0}^{N_{\phi}}\nu_{i}(i{-}N_{\phi}/2){=}N_{e}/2$. The two conditions are equivalent to the constraints that a single quasihole has a charge $e_{1{-}\text{qh}}{=}{-}1/m$ and a dipole moment $d_{1{-}\text{qh}}{=}(1{-}1/m)/2$ in the thermodynamic limit. Both of these relations are topological and the focus of $K$-matrix Luttinger liquid theory, thus robust against perturbations [19]. From finite-size-scaling, we find the following values of the parameters in the thermodynamic limit [see Appendix. B for details]:
@@
III Stacked Laughlin quasiholes
The damped oscillation model can also be applied to $n$-stacked quasiholes created by inserting $n$ fluxes at the same location. The real-space density of such quasiholes gets increasingly complex with increasing $n$. Even for IQH fluid, the $n$-stacked holes have $\delta\rho(r){=}{-}1/(2\pi)\exp\left({-}r^{2}/2\right)(\sum_{i{=}0}^{n{-}1}r^{2i}/2^{i}i!)$, which is no longer a simple Gaussian distribution for $n{>}1$. Focusing on the FQH system with multiple quasiholes, we again look at its occupation-number density which reveals interesting physics and is no more complex even for $n{>}1$. The values of the characteristic lengths $k_{1}^{(n)}$ and $\lambda_{1}^{(n)}$ for different values of $n$ are shown in Figs. 3(a) and 3(b), where both $k_{1}^{(n)}$ and $\lambda_{1}^{(n)}$ increase as $n$ gets larger. Furthermore, the complicated real-space densities can be easily restored with the single-particle wave functions.
@@ III Stacked Laughlin quasiholes
The $m$-stacked quasihole at $\nu{=}1/m$ deserves special attention because its density distribution is proportional to the pair-correlation function of the Laughlin ground state. The density distribution of $m$-Laughlin quasiholes is $\rho(r){=}\mathcal{N}_{1}\int\prod_{i{=}2}^{N_{e}^{\prime}}d\bm{r}_{i}|\Psi_{L}^{m{-}\text{qh}}(\bm{r},\bm{r}_{2},{\ldots},\bm{r}_{N_{e}^{\prime}})|^{2},$ where $\mathcal{N}_{1}$ is a constant; and the pair-correlation function is $g(r){=}\mathcal{N}_{2}\int\prod_{i{=}3}^{N_{e}}d\bm{r}_{i}|\Psi_{L}(0,\bm{r},\bm{r}_{3},{\ldots},\bm{r}_{N_{e}})|^{2}.$ By substituting the explicit wave function $\Psi_{L}$ and $\Psi_{L}^{m{-}\text{qh}}$ into the expression of $g(r)$ and $\rho(r)$, respectively, one can show that $\rho(r)$ is proportional to $g(r)$ when $N_{e}{=}N_{e}^{\prime}{+}1$. Therefore, the decay length of the density of $m$-stacked Laughlin quasiholes is equal to the correlation length $\lambda_{\text{cor}}$ between electrons:
@@
Our result is thus useful for the large-scale numerical computation of ground-state variational energies. One can use our modeled $g(r)$ to calculate the per-particle variational energy of the Laughlin state in terms of any general interaction $V(r)$ through the formula $V{=}(\bar{\rho}/2)\int d^{2}\,\bm{r}V(r)[g(r){-}1]$ [31], where $\bar{\rho}$ is the average density. For the Coulomb interaction $V(r){=}1/r$, the calculated energy $V$ with our modeled $g(r)$ is $-0.4096$ for $\nu{=}1/3$ and $-0.3278$ for $\nu{=}1/5$ [see Appendices. C and E], which are very close to the conjectured thermodynamic values $-0.4098$ [32, 33] and $-0.3275(1)$ [34] obtained from the extrapolation of small-size exact diagonalization results and large-scale Monte Carlo calculations. Thus for other realistic interactions, the thermodynamic variational energies can now be computed very efficiently with our modeled $g(r)$ [see Appendices. C and E]. Besides, our method does not suffer from the systemic error of the polynomial expansion method at small $r$ [29].
In the limit of $n{\rightarrow}\infty$, the corresponding “Laughlin quasihole” also deserves special attention as it represents the edge of the FQH fluid. Its density profile can also be well-fitted with a damped model in the occupation-number space with the characteristic lengths being $k_{1}^{\text{edge}}{=}1.49$ and $\lambda_{1}^{\text{edge}}{=}1.51$ [see Appendix. D] which are just the limiting values of $n$-quasiholes as shown in Fig. 3.
@@
Given the near-perfect fitting of such a simple model with only two fitting parameters [see Appendices. B-F], we conjecture that the damped oscillation model captures the main features of the universal oscillations in the Laughlin FQH fluid, and in principle can be derived. While we are not able to accomplish that here, a phenomenological model for Laughlin quasiholes at general fillings can be proposed based on these results. We view the system as a damped oscillator with wave number $k^{\prime}_{1}{=}\sqrt{k_{1}^{2}{+}\lambda_{1}^{-2}}$ serving as the intrinsic “wave number” which is a little larger than the wave number of the Wigner crystal $k_{\text{Wigner}}$ for small $m$ [see Fig. 3(c)]. The damping results from a “frictional force” which is proportional to the density $\nu{=}1/m$. The differential equation that describes this system is
@@
where $\mathcal{O}(\delta n^{2})$ denote terms that can potentially arise from non-linear effects which we ignore here. The linear solution of Eq. (4) is just our damped model given in Eq. (6), which is known to model the classical two-dimensional shallow-water wave near a specified moving boundary [36, 37]. Note that the model proposed in Eq. (4) is different from that proposed in Eq. (23) of Ref. [21], which does not consider the damped term and assumes a different intrinsic wave number at the level of linear response. Deriving this simple model from the microscopic wave function and capturing the non-linear effects remains an open problem [21, 22].
IV Quasiholes for realistic interactions
Going beyond the model Laughlin quasihole wave functions, the density profile of quasiholes from realistic interactions becomes more complicated, because such quasiholes are dressed by neutral excitations [38, 39]. Here we study the quasihole in the presence of the following tunable interaction